1)                 If f(x) = 3x2 – 2x + 5, find f[1, 2], f[2, 3] and f[1, 2, 3].


1
Expert's answer
2021-07-21T15:01:05-0400

(1.) The function "f(x)=3x^2-2x+5" is continuous for all "x\in \R" as polynomial.

and the first derivative of this function f(x) is



"f'(x)=6x-2""f'(x)=0\implies 6x-2=0\implies x=\dfrac{1}{3}"

If "x>\dfrac{1}{3}, f'(x)>0, f(x)" is strictly increasing.




"f(1)=3(1)^2-2(1)+5=6""f(2)=3(2)^2-2(2)+5=13""f(3)=3(3)^2-2(3)+5=26""\implies f:[1,2]\to[6,13]""\implies f:[2,3]\to[13,26]""\implies f:\{1,2,3\}\to\{6, 13,26\}"